OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

I'm so confused

OpenStudy (campbell_st):

ok.... so looking at the 2 equations..... the easiest letter to eliminate is y. they have opposite sides... x - y = 19 adding the 2 equations will eliminate y and allow you to 2x + y = 2 solve for x ----------- 3x = 21 now divide both sides of the equation to find x. x = 7 to find the value of y, substitute x = 7 into either original equation I'll choose the 1st equation 7 - y = 19 subtract 7 from both sides of the equation -y = 12 divide by -1 to get the value of y y = -12 so the solution to these 2 equations is x = 7 and y = -12 All the solution means that if you graphed these to lines on the number planes you would see they intersect at (7, -12) hope that makes sense.

OpenStudy (campbell_st):

other methods you can use to solve these equations are 1. graphing... but given the solution (7, -12) you would need a large number plane. 2. Substitution method. which means taking an equation, changing the subject and then substituting into the other equation. If I choose the 1st equation and make x the subject you get x = y + 19 now substitute it into the 2nd equation 2(y + 19) + y = 2 distributing 2y + 38 + y = 2 collect like terms 3y + 38 = 2 and if you continue you'll find y = -12... same as before... and substituting you would get x = 7 For this question elimination is the easier method because the y terms have the same coefficient and opposite signs... so adding them well cause them to be eliminated. Hope it all makes sense

OpenStudy (anonymous):

Thx much better

OpenStudy (campbell_st):

glad to help